Solutions of sodium hydroxide (volume = 90.0 mL) and nitric acid (volume = 95.0 mL) were mixed together in a neutralization reaction.

Calculate the mass of the combined solution.

We can assume the density of the solution to be the same as water, 1.00 g/cm3.


Give you answer to three significant figures.



1
Expert's answer
2020-08-26T03:11:02-0400

Volume of Sodium Hydroxide "=90.0\ ml"

Volume of Nitric Acid"=95.0 \ ml"

Total Volume "=(90.0+95.0)\ ml=185\ ml"

Density of solution "= 1.00\ g/cm^3=1.00\ g/ml\ (\because 1 cm^3=1\ ml)"

Mass of solution will be "= V\times\rho=185\times 1.00 \ g= 185.00\ g"

"=185 \ g" (In three significant figures)